In Mathematics: how much interest will you have to pay for a credit card balance of $554 that is one month overdue if a 20% annual rate is charged?
Answer #1
0.2/12 = 0.01666666 interest rate per month

554 * 0.0166666 = 9.2332964
554 + 9.23 = 563.23

563.23 * 0.01666666 = 9.39

9.39 + 9.23 = ? interest charges
answered by: lili